41. Jack can get a pail of water in 5 minutes. Jill can get a pail of water in 3 minutes. If they work together (and don’t fall down) how long will it take to get a pail of water?

Answer by checkley75(3666)

5*3/(5+3)=15/8=1.875 hours to fill a pail if they work together.